# Gaussian Process Upper Confidence Bound Achieves Nearly-Optimal Regret in Noise-Free Gaussian Process Bandits

### v1

@@ 3 Refined Regret Upper Bound of Noise-Free GP-UCB
Under the fixed $L$, $d$, $\ell$, and $\nu$, the above lemma claims
$\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})=O\left(\sqrt{T}\exp\left(-\frac{1}{2}\widetilde{C}_{\mathrm{SE}}T^{\frac{1}{d+1}}\right)\right),~{}\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})=O(1),~{}\mathrm{and}$ (15)
$\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})=O\left(T^{-\frac{\nu}{d}}(\ln T)^{\frac{\nu}{d}}\right),~{}\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})=\begin{cases}O\left(T^{\frac{d-\nu}{d}}(\ln T)^{\frac{\nu}{d}}\right)~{}~{}&\mathrm{if}~{}~{}d>\nu,\\ O\left((\ln T)^{2}\right)~{}~{}&\mathrm{if}~{}~{}d=\nu,\\ O(1)~{}~{}&\mathrm{if}~{}~{}d<\nu,\end{cases}$ (16)
for $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$, respectively. Combining the above equations with Eqs. (9) and (10), we obtain the desired results.
Remark 1 (Generality of Lemma 3).
We would like to highlight that Lemma 3 always holds in any input sequence, in contrast to the existing algorithm-specific upper bounds (Salgia et al., 2024; Iwazaki and Takeno, 2025). Since the existing noisy GP bandits theory often leverage the upper bound of $\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ or $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ from (Srinivas et al., 2010), we expect that the various existing theory of the noisy setting algorithms can be extended to the corresponding noise-free setting by directly replacing the existing noisy upper bounds of (Srinivas et al., 2010) with Lemma 3. For example, the analysis for GP-TS (Chowdhury and Gopalan, 2017), GP-UCB and GP-TS under Bayesian setting (Srinivas et al., 2010; Russo and Van Roy, 2014), contextual setting (Krause and Ong, 2011), GP-based level-set estimation (Gotovos et al., 2013), multi-objective setting (Zuluaga et al., 2016), robust formulation (Bogunovic et al., 2018), and so on.
3.1 Proof Sketch of Lemma 3
In this section, we describe the proof sketch of Lemma 3, while we give its full proof in Appendix A.2. We consider the following more general lemmas from which Lemma 3 follows.
With some elemental calculations, we can confirm that the condition $\forall t\geq\overline{T},\gamma_{t}(\lambda_{t}^{2})\leq(t-1)/3$ of the above lemmas holds with $\lambda_{t}^{2}=t\exp(-\widetilde{C}_{\mathrm{SE}}t^{\frac{1}{d+1}}),\overline{T}=\overline{T}_{\mathrm{SE}}$ and $\lambda_{t}^{2}=\widetilde{C}_{\mathrm{Mat}}t^{-\frac{2\nu}{d}}(\ln t)^{\frac{2\nu}{d}},\overline{T}=\overline{T}_{\mathrm{Mat}}$ for $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$, respectively. See Appendix A.2 for details. Lemma 3 follows from the aforementioned setting of $\lambda_{t}^{2}$, $\overline{T}$, and Lemmas 4 and 5.
@@ 3.1 Proof Sketch of Lemma 3
Instead of directly treating the noiseless posterior standard deviation, we study its upper bound with the posterior standard deviation of some noisy GP model. Here, let us denote $\sigma_{\lambda^{2}}^{2}(\bm{x};\mathbf{X}_{t-1})$ as the posterior variance under the noisy GP-model with the strictly positive variance parameter $\lambda^{2}>0$, which is defined as
Lemma 6 (Elliptical potential count lemma, Lemma D.9 in (Flynn and Reeb, 2024) or Lemma 3.3 in (Iwazaki and Takeno, 2025)).
Fix any $T\in N_{+}$, any sequence $\bm{x}_{1},\ldots,\bm{x}_{T}\in\mathcal{X}$, and $\lambda>0$. Define $\mathcal{T}$ as $\mathcal{T}=\{t\in[T]\mid\lambda^{-1}\sigma_{\lambda^{2}}(\bm{x}_{t};\mathbf{X}_{t-1})>1\}$, where $\mathbf{X}_{t-1}=(\bm{x}_{1},\ldots,\bm{x}_{t-1})$. Then, the number of elements of $\mathcal{T}$ satisfies $|\mathcal{T}|\leq 3\gamma_{T}(\lambda^{2})$.
Proof of Lemma 5.
If $T<\overline{T}$, Eq. (17) is clearly holds from the assumption $\forall\bm{x}\in\mathcal{X},k(\bm{x},\bm{x})\leq 1$. Hereafter, we focus on $T\geq\overline{T}$. First, by following the same argument of Lemma 4, we can confirm that there exists the index $\widetilde{t}_{T}\leq T$ such that $\sigma(\bm{x}_{\widetilde{t}_{T}};\mathbf{X}_{\widetilde{t}_{T}-1})\leq\sigma_{\lambda_{T}^{2}}(\bm{x}_{\widetilde{t}_{T}};\mathbf{X}_{\widetilde{t}_{T}-1})\leq\lambda_{T}$. Here, we define the sequence $(\bm{x}_{t}^{(T-1)})_{t\in[T-1]}$ as the sequence that $\bm{x}_{\widetilde{t}}$ is eliminated from $(\bm{x}_{t})_{t\in[T]}$; namely, we set $\bm{x}_{t}^{(T-1)}=\mbox{1}\mbox{l}\{t<\widetilde{t}_{T}\}\bm{x}_{t}+\mbox{1}\mbox{l}\{t\geq\widetilde{t}_{T}\}\bm{x}_{t+1}$ for any $t\in[T-1]$. Furthermore, we define $\mathbf{X}_{t}^{(T-1)}=(\bm{x}_{1}^{(T-1)},\ldots,\bm{x}_{t}^{(T-1)})$. Under these notations, we have
Then, we observe that there exists the index $\widetilde{t}_{T-1}\leq T-1$ such that $\sigma(\bm{x}_{\widetilde{t}_{T-1}}^{(T-1)};\mathbf{X}_{\widetilde{t}_{T-1}-1}^{(T-1)})\leq\sigma_{\lambda_{T-1}^{2}}(\bm{x}_{\widetilde{t}_{T-1}}^{(T-1)};\mathbf{X}_{\widetilde{t}_{T-1}-1}^{(T-1)})\leq\lambda_{T-1}$ by the application of Lemma 6 for the new sequence $(\bm{x}_{t}^{(T-1)})$. Again, by setting $\bm{x}_{t}^{(T-2)}=\mbox{1}\mbox{l}\{t<\widetilde{t}_{T-1}\}\bm{x}_{t}^{(T-1)}+\mbox{1}\mbox{l}\{t\geq\widetilde{t}_{T-1}\}\bm{x}_{t+1}^{(T-1)}$ and $\mathbf{X}_{t}^{(T-2)}=(\bm{x}_{1}^{(T-2)},\ldots,\bm{x}_{t}^{(T-2)})$ for any $t\in[T-2]$, we have
$\sum_{t\in[T-1]}\sigma\left(\bm{x}_{t}^{(T-1)};\mathbf{X}_{t-1}^{(T-1)}\right)\leq\sum_{t\in[T-1]\setminus\{\widetilde{t}_{T-1}\}}\sigma\left(\bm{x}_{t}^{(T-1)};\mathbf{X}_{t-1}^{(T-1)}\right)+\lambda_{T-1}\leq\sum_{t\in[T-2]}\sigma\left(\bm{x}_{t}^{(T-2)};\mathbf{X}_{t-1}^{(T-2)}\right)+\lambda_{T-1}.$ (20)
$\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})\leq\sum_{t=1}^{\overline{T}-1}\sigma\left(\bm{x}_{t}^{(\overline{T}-1)};\mathbf{X}_{t-1}^{(\overline{T}-1)}\right)+\sum_{t=\overline{T}}^{T}\lambda_{t}\leq\overline{T}-1+\sum_{t=\overline{T}}^{T}\lambda_{t}.$ (21)
4 Lower Bound for Cumulative Regret under Noise-Free Setting
Our analysis shows the GP-UCB achieves nearly optimal cumulative regret for $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$ with $d\leq\nu$ since the constant $O(1)$ cumulative regret is unavoidable. The remaining question is whether GP-UCB is nearly optimal or not under $d>\nu$. Our following result gives the affirmative answer by showing the lower bound for the deterministic regret.
Theorem 7 (Lower bound for deterministic regret under $k=k_{\mathrm{Mat\acute{e}rn}}$).
Suppose $\mathcal{X}=[0,1]^{d}$, $k=k_{\mathrm{Mat\acute{e}rn}}$, and $d$, $\ell$, and $\nu$ are fixed constants with $d>\nu$. Then, for any algorithm (including both deterministic and stochastic algorithms) and $T\geq 2$, there exists a function $f$ and an event [^3] We can take the event whose probability of occurrence is at least $1/T$ when the algorithm is stochastic. such that $\|f\|_{k}\leq B$ and $R_{T}\geq CBT^{\frac{d-\nu}{d}}$. Here, $C>0$ is the constant that may depend on $d$, $\ell$, and $\nu$.
We would like to emphasize that the above result only shows the lower bound for the deterministic regret; namely, there exists the possibility for achieving smaller expected regret $\mathbb{E}[R_{T}]$ than $\Omega(BT^{(d-\nu)/d})$ by the randomness of the algorithm. We leave the analysis for the expected regret lower bound for future research; however, note that our deterministic regret lower bound is enough to guarantee the near optimality of our analysis of GP-UCB.
Proof sketch.
As with the proof of noisy lower bound (Scarlett et al., 2017; Cai and Scarlett, 2021), we consider the finite collection of bump functions, whose unique maximizers are different and the function values are 00 on the neighborhood of the other function maximizers. For any small $\epsilon>0$, Cai and Scarlett (2021) shows that we can construct $M=\Theta\left((B/\epsilon)^{d/\nu}\right)$ distinct functions whose maximum and RKHS norm upper bounds are $\Theta(\epsilon)$ and $B$, respectively. Here, let us fix any input sequence $\bm{x}_{1},\ldots,\bm{x}_{M-1}\in\mathcal{X}$ generated by the algorithm when the underlying function is $f=0$. Then, the same input sequence is also generated by the algorithm when the underlying function is the bump function, whose function values are exactly zero at $\bm{x}_{1},\ldots,\bm{x}_{M-1}$. Furthermore, under such bump function, the algorithm suffers from at least $\Omega((M-1)\epsilon)=\Omega((T-1)\epsilon)=\Omega(T\epsilon)$ regret when $\bm{x}_{1},\ldots,\bm{x}_{M-1}$ is generated and $M=T$. Since the condition $M=T$ implies $\epsilon=\Theta(BT^{-\frac{\nu}{d}})$, we obtain the desired lower bound $\Omega(BT^{(d-\nu)/d})=\Omega(T\epsilon)$.
5 Conclusion
This paper shows that the GP-UCB achieves nearly optimal regret by proving the new regret upper and lower bound for noise-free GP bandits. The key theoretical component of our analysis is the tight upper bound of the posterior standard deviations of GP tailored to a noise-free setting (Lemma 3). As remarked in Section 3, Lemma 3 can be applicable beyond the analysis of GP-UCB. Specifically, we expect that the various existing theoretical results for noisy GP bandits settings can translate into its noise-free setting by replacing the existing noisy upper bound of the posterior standard deviations with Lemma 3. For this reason, we believe that our result marks an important step for the future development of noise-free GP bandit algorithms.
Appendix A Proofs for Section 3
A.1 Proof of Theorems 1 and 2
Lemma 8 (Deterministic confidence bound for noise-free setting, Lemma 11 in [Lyu et al., 2019] or Proposition 1 in [Vakili et al., 2021a]).
where $\mathbf{X}_{t}=(\bm{x}_{1},\ldots,\bm{x}_{t})$.
Although the remaining parts of the proofs are the well-known results of GP-UCB, we give the details for completeness. Based on the above lemma, we show Eqs. (9) and (10). Regarding $R_{T}$, we have
$R_{T}$ $=\sum_{t=1}^{T}f(\bm{x}^{\ast})-f(\bm{x}_{t})$ (23)
$=2B\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t}),$ (26)
where the first inequality follows from Lemma 8, and the second inequality follows from the UCB-selection rule of $\bm{x}_{t}$. Similarly to the cumulative regret, we have
$r_{T}$ $=f(\bm{x}^{\ast})-f(\widehat{\bm{x}}_{T})$ (27)
$\leq\min_{t\in[T]}f(\bm{x}^{\ast})-f(\bm{x}_{t})$ (28)
$=2B\min_{t\in[T]}\sigma(\bm{x}_{t};\mathbf{X}_{t}),$ (31)
where the first inequality follows from the definition of $\widehat{\bm{x}}_{T}$. Finally, the desired results are obtained by combining the above inequalities with Eqs. (15) and (16). ∎
A.2 Proof of Lemma 3
@@
$\Leftrightarrow\widetilde{C}_{\mathrm{SE}}\leq\left(\frac{1}{6C_{\mathrm{SE}}}\right)^{d+1},$ (37)
where the second line follows from the inequality $t-1\geq t/2$ for all $t\geq\overline{T}_{\mathrm{SE}}\geq 2$. By noting the definition of $\widetilde{C}_{\mathrm{SE}}$, we conclude that $\forall t\geq\overline{T}_{\mathrm{SE}},\gamma_{t}(\lambda_{t}^{2})\leq C_{\mathrm{SE}}\widetilde{C}_{\mathrm{SE}}^{d+1}t\leq\frac{t-1}{3}$ from the above inequalities, which implies Lemmas 4 and 5 holds with $\lambda_{t}^{2}=t\exp(-\widetilde{C}_{\mathrm{SE}}t^{\frac{1}{d+1}})$ and $\overline{T}=\overline{T}_{\mathrm{SE}}$. Eq. (11) directly follows from Lemma 4 with the fact $\sigma(\bm{x}_{t};\mathbf{X}_{t-1})\leq k(\bm{x}_{t},\bm{x}_{t})\leq 1$. As for Eq. (12), Lemma 5 implies
@@
where the second line follows from the fact that the function $g(t)\coloneqq t\exp(-\widetilde{C}_{\mathrm{SE}}t^{1/(d+1)})$ is non-increasing for $t\geq\overline{T}_{\mathrm{SE}}-1$. Indeed, we have
@@
which implies $g^{\prime}(t)\leq 0$ for $t\geq\overline{T}_{\mathrm{SE}}-1\geq(d+1)^{d+1}/\widetilde{C}_{\mathrm{SE}}^{d+1}$. Regarding the quantity $\int_{1}^{T}\sqrt{t}\exp\left(-\frac{1}{2}\widetilde{C}_{\mathrm{SE}}t^{\frac{1}{d+1}}\right)\text{d}t$, we further obtain the following upper bound with $C\coloneqq\widetilde{C}_{\mathrm{SE}}/2>0$:
@@
Next, when $k=k_{\mathrm{Mat\acute{e}rn}}$, we set $\lambda_{t}^{2}=\widetilde{C}_{\mathrm{Mat}}t^{-\frac{2\nu}{d}}(\ln t)^{\frac{2\nu}{d}}$ and $\overline{T}=\max\{4,\underline{T}_{\mathrm{Mat}},\underline{T}_{\mathrm{Mat}}^{(\lambda)}\}$ with $\widetilde{C}_{\mathrm{Mat}}=\left(2+\frac{2\nu}{d}\right)^{\frac{2\nu}{d}}(6C_{\mathrm{Mat}})^{1+\frac{2\nu}{d}}$. Then, for any $t\geq\overline{T}_{\mathrm{Mat}}$, we have
@@
$\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ $\leq\overline{T}_{\mathrm{Mat}}-1+\sum_{t=\overline{T}_{\mathrm{Mat}}}^{T}\lambda_{t}$ (54)
@@
where the last line follows from the standard property of Gamma function: $\int_{0}^{\infty}e^{-\lambda u}u^{b}\text{d}u=\Gamma(b+1)/\lambda^{b+1}$ for any $\lambda>0$ and $b>-1$ (e.g., Equation 6.1.1 in [Abramowitz and Stegun, 1968]). ∎
Appendix B Proof of Theorem 7
We leverage the following lemma to define the base function, which is used to construct a finite collection of “hard” functions.
Lemma 9 (Lemma 4 in [Cai and Scarlett, 2021]).
Fix any $k=k_{\mathrm{Mat\acute{e}rn}}$, $\epsilon>0$, and $\omega>0$. Let $h(\bm{x})\coloneqq\exp\left(-1/(1-\|\bm{x}\|_{2}^{2})\right)\mbox{1}\mbox{l}\{\|\bm{x}\|_{2}<1\}$ be the $d$-dimensional bump function, and define $g(\bm{x})=\frac{2\epsilon}{h(\bm{0})}h(\bm{x}/\omega)$. Then, the function $g$ satisfies the following properties:
  1. 1.
    $g(\bm{x})=0$ for all $\bm{x}$ outside the $\ell_{2}$-ball of radius $\omega$ centered at the origin.
  2. 2.
    $g(\bm{x})\in[0,2\epsilon]$ for all $\bm{x}$ and $g(\bm{0})=2\epsilon$.
  3. 3.
    $\|g\|_{k}\leq\frac{2c_{1}\epsilon}{h(0)}(1/\omega)^{\nu}\|h\|_{k}$, where $c_{1}$ is some finite constant. In particular, we have $\|g\|_{k}\leq B$ when $\omega=(2c_{1}\epsilon\|h\|_{k}/(h(\bm{0})B))^{1/\nu}$.
Note that $k=k_{\mathrm{Mat\acute{e}rn}}$ is stationary, any shifted function $f(\bm{x})=g(\bm{x}-\bm{a})$ maintain the RKHS norm (e.g., Lemma 1 in [Scarlett et al., 2017]). Therefore, given some $\epsilon>0$, by shifting the above function $g$ with $\omega\coloneqq\omega_{\epsilon}=(2c_{1}\epsilon\|h\|_{k}/(h(\bm{0})B))^{1/\nu}$, we can construct the function class $\mathcal{F}_{\epsilon}=\{f_{1}^{(\epsilon)},\ldots,f_{M_{\epsilon}}^{(\epsilon)}\}$ such that $\|f_{i}^{(\epsilon)}\|_{k}\leq B$ and each function in $\mathcal{F}_{\epsilon}$ has unique non-zero support. The maximum size $M_{\epsilon}>0$ of $\mathcal{F}_{\epsilon}$ for such construction is obtained by the $\omega_{\epsilon}$-packing number of $\mathcal{X}\coloneqq[0,1]^{d}$. As with the proof of existing lower bound [Scarlett et al., 2017, Cai and Scarlett, 2021], we set $M_{\epsilon}$ as
$M_{\epsilon}=\left\lfloor\frac{1}{\omega_{\epsilon}}\right\rfloor^{d}=\left\lfloor\left(\frac{h(\bm{0})B}{6\epsilon c_{1}\|h\|_{k}}\right)^{1/\nu}\right\rfloor^{d},$ (63)
which is clearly less than $\omega_{\epsilon}$-packing number of $\mathcal{X}\coloneqq[0,1]^{d}$. Furthermore, since each function in $\mathcal{F}_{\epsilon}$ has unique non-zero support, we can divide $\mathcal{X}$ into the $M_{\epsilon}$ disjoint regions $(\mathcal{R}_{i}^{(\epsilon)})_{i\in[M_{\epsilon}]}$ based on $\mathcal{F}_{\epsilon}$ such that $\mathcal{X}=\bigsqcup_{i\in[M_{\epsilon}]}\mathcal{R}_{i}^{(\epsilon)}$ and $\{\bm{x}\in\mathcal{X}\mid f_{i}^{(\epsilon)}(\bm{x})>0\}\subset\mathcal{R}_{i}^{(\epsilon)}$. Here, we fix $\epsilon$ as $\epsilon=\left(\frac{h(\bm{0})}{6c_{1}\|h\|_{k}}\right)BT^{-\frac{\nu}{d}}$, which implies $M_{\epsilon}=T$. Then, for a given algorithm $\pi$, we pick the region $\mathcal{R}_{j}^{(\epsilon)}$ such that the algorithm query sequence $\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}$ under $f\coloneqq f_{0}=0$ satisfies
$\mathbb{P}_{f_{0},\pi}\left(\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\notin\mathcal{R}_{j}^{(\epsilon)}\right)=\max_{i\in[M_{\epsilon}]}\mathbb{P}_{f_{0},\pi}\left(\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\notin\mathcal{R}_{i}^{(\epsilon)}\right).$ (64)
Note that such $\mathcal{R}_{j}^{(\epsilon)}$ also satisfies $\mathbb{P}_{f_{0},\pi}\left(\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\notin\mathcal{R}_{j}^{(\epsilon)}\right)\geq 1/M_{\epsilon}$, since any realized sequence $\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}$ has at least one region in which it is not contained due to $M_{\epsilon}=T$; namely, $\sum_{i=1}^{M_{\epsilon}}\mathbb{P}_{f_{0},\pi}\left(\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\notin\mathcal{R}_{i}^{(\epsilon)}\right)=\sum_{i=1}^{M_{\epsilon}}\mathbb{P}_{f_{0},\pi}\left(\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\in\mathcal{X}\setminus\mathcal{R}_{i}^{(\epsilon)}\right)=1$. Furthermore, from the definition of $\mathcal{R}_{i}^{(\epsilon)}$, $\mathbb{P}_{f_{j},\pi}\left(\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\notin\mathcal{R}_{j}^{(\epsilon)}\right)=\mathbb{P}_{f_{0},\pi}\left(\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\notin\mathcal{R}_{j}^{(\epsilon)}\right)\geq 1/M_{\epsilon}$. By noting the sequence $\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\notin\mathcal{R}_{j}^{(\epsilon)}$ suffer from at least $2\epsilon(M_{\epsilon}-1)$ regret under $f=f_{j}$, we have the following:
$\mathbb{P}_{f_{j},\pi}\left(R_{T}\geq 2\epsilon(M_{\epsilon}-1)\right)\geq 1/M_{\epsilon}=1/T.$ (65)
Therefore, for any $T\geq 2$ and any algorithm $\pi$, there exists $f\in\mathcal{H}_{k}$ such that $\|f\|_{k}\leq B$ and the following inequality holds with probability at least $1/T$.
$R_{T}\geq 2\epsilon(M_{\epsilon}-1)=2\left(\frac{h(\bm{0})}{6c_{1}\|h\|_{k}}\right)BT^{-\frac{\nu}{d}}(T-1)\geq\left(\frac{h(\bm{0})}{6c_{1}\|h\|_{k}}\right)BT^{\frac{d-\nu}{d}}.$ (66)
By setting $C=\left(\frac{h(\bm{0})}{6c_{1}\|h\|_{k}}\right)$ in the above inequality, we obtain the desired statement. ∎
Remark 2.
The above proof does not lead to the same order of expected regret as the deterministic regret lower bound in Theorem 7, since the probability of the event, which the lower bound holds with, may decay in the order of $1/T$. Therefore, the only thing we can claim is $\mathbb{E}_{f_{j}}[R_{T}]\geq\mathbb{E}_{f_{j}}[\mbox{1}\mbox{l}\{\bm{x}_{1},\ldots,\bm{x}_{M_{\epsilon}-1}\notin\mathcal{R}_{j}^{(\epsilon)}\}R_{T}]\geq\Omega(BT^{-\nu/d})$, which is worse than trivial constant regret lower bound $\Omega(1)$.

### v2

@@ A.2 Proof of Lemma 3
The upper bound in Lemma 8 depends on the quantities $\widetilde{C}_{\mathrm{SE}}$, $\overline{T}_{\mathrm{SE}}$, $\widetilde{C}_{\mathrm{Mat}}$, $\overline{T}_{\mathrm{Mat}}>0$, which are related to the implied constants in the upper bounds of the MIG. However, under the fixed $d$, $\ell$, and $\nu$, $\widetilde{C}_{\mathrm{SE}}$, $\overline{T}_{\mathrm{SE}}$, $\widetilde{C}_{\mathrm{Mat}}$, $\overline{T}_{\mathrm{Mat}}>0$ are also constants, which implies the conclusions of Lemma 3.
@@
$\Leftrightarrow\widetilde{C}_{\mathrm{SE}}\leq\left(\frac{1}{6C_{\mathrm{SE}}}\right)^{\frac{1}{d+1}},$ (40)
where the second line follows from the inequality $t-1\geq t/2$ for all $t\geq\overline{T}_{\mathrm{SE}}\geq 2$. By noting the definition of $\widetilde{C}_{\mathrm{SE}}$, we conclude that $\forall t\geq\overline{T}_{\mathrm{SE}},\gamma_{t}(\lambda_{t}^{2})\leq C_{\mathrm{SE}}\widetilde{C}_{\mathrm{SE}}^{d+1}t\leq\frac{t-1}{3}$ from the above inequalities, which implies that Lemmas 4 and 5 hold with $\lambda_{t}^{2}=t\exp(-\widetilde{C}_{\mathrm{SE}}t^{\frac{1}{d+1}})$ and $\overline{T}=\overline{T}_{\mathrm{SE}}$. Eq. (31) directly follows from Lemma 4 using the fact $\sigma(\bm{x}_{t};\mathbf{X}_{t-1})\leq k(\bm{x}_{t},\bm{x}_{t})\leq 1$. As for Eq. (32), Lemma 5 implies
@@
where the second line follows from the fact that the function $g(t)\coloneqq t\exp(-\widetilde{C}_{\mathrm{SE}}t^{1/(d+1)})$ is non-increasing for $t\geq\overline{T}_{\mathrm{SE}}-1$. In fact, we have
@@
which implies $g^{\prime}(t)\leq 0$ for $t\geq\overline{T}_{\mathrm{SE}}-1\geq(d+1)^{d+1}/\widetilde{C}_{\mathrm{SE}}^{d+1}$. To bound the quantity $\int_{1}^{T}\sqrt{t}\exp\left(-\frac{1}{2}\widetilde{C}_{\mathrm{SE}}t^{\frac{1}{d+1}}\right)\text{d}t$, we further derive the following upper bound with $C\coloneqq\widetilde{C}_{\mathrm{SE}}/2>0$:
@@
Next, when $k=k_{\mathrm{Mat\acute{e}rn}}$, we set $\lambda_{t}^{2}=\widetilde{C}_{\mathrm{Mat}}t^{-\frac{2\nu}{d}}(\ln t)^{\frac{2\nu}{d}}$ and $\overline{T}=\max\{4,\underline{T}_{\mathrm{Mat}},\underline{T}_{\mathrm{Mat}}^{(\lambda)}\}$ with $\widetilde{C}_{\mathrm{Mat}}=\left(2+\frac{2\nu}{d}\right)^{\frac{2\nu}{d}}(6C_{\mathrm{Mat}})^{1+\frac{2\nu}{d}}$. Then, for any $t\geq\overline{T}_{\mathrm{Mat}}$, it holds that
@@
$\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ $\leq\overline{T}_{\mathrm{Mat}}+\sum_{t=\overline{T}_{\mathrm{Mat}}}^{T}\lambda_{t}$ (57)
@@
where the last line follows from the standard property of Gamma function: $\int_{0}^{\infty}e^{-\lambda u}u^{b}\text{d}u=\Gamma(b+1)/\lambda^{b+1}$ for any $\lambda>0$ and b>1b>-1 (e.g., Equation 6.1.1 in [abramowitz1968handbook]). ∎
Appendix B Detail of the Experiment
B.1 Experimental settings for Figure 1
We give the detailed experimental settings used to plot Figure 1.
  • Objective function. We define the true underlying objective function as $f(\cdot)=\sum_{m=1}^{50}c_{m}k(\bm{x}^{(m)},\cdot)$, where $c_{m}\sim\mathrm{Uniform}([-1,1])$ and $\bm{x}^{(m)}\sim\mathrm{Uniform}([0,1]^{2})$ are independently generated random variables. Note that by the definition of the RKHS, $f\in\mathcal{H}_{k}$ with $\|f\|_{k}=\sqrt{\sum_{m=1}^{50}\sum_{\widetilde{m}=1}^{50}c_{m}c_{\widetilde{m}}k(\bm{x}^{(m)},\bm{x}^{(\widetilde{m})})}$.
  • Kernel. We fix the lengthscale parameter $\ell$ as $\ell=0.25$ in all experiments. We use the same kernel function in the GP model used in the algorithms as the kernel leveraged to generate the objective function.
  • Other parameters. We define the input domain $\mathcal{X}$ as the uniformly aligned $50\times 50$ grid points on $0,1]^{2}$. In all algorithms, we set the confidence width parameter $\beta^{1/2}$ to the exact RKHS norm. Furthermore, we set the initial batch size in PE and REDS as $5$. Finally, we define the common initial point $\bm{x}_{1}$ for all the algorithms as the uniformly sampled point from $\mathcal{X}$.
In the above-described setting, we conduct experiments with $3000$ different seeds.
B.2 Comparison between EI and GP-UCB
Under the same setting as the previous subsection, we also compare GP-UCB’s empirical performance with that of EI in simple regret minimization. Figure 2 shows the results. We can confirm that, although GP-UCB achieves nearly-optimal worst-case regret, the empirical performance of GP-UCB is consistently worse than that of EI.
Figure 2: Comparison between GP-UCB and EI in simple regret minimization over $100$ different seeds. We conduct experiments with Matérn kernel under $\nu=3/2$ (left) and $\nu=5/2$ (right).